Solve the following equation by elimination.
x+y = 500
0.09x+20 = .11y
No clue where to start please help.

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

IN ORDER TO ELIMINATE ONE VARIABLE WE NEED TO MULTIPLY THE FIRST EQUATION BY
-.09 TO ELIMINATE THE X VARIABLE THUS
-.09X+-09Y=-45
.09X-.11Y=-20 NOW WE CAN ADD THESE TWO EQUATION
-------------------------------
-.2Y=-65 OR Y=-65/-.2 OR Y=325 & X+325=500 OR X=500-325 OR X=175
PROOF .09*175+20=.11*325 OR 15.75+20=35.75 OR 35.75=35.75

Answer by stanbon(48558)

1st: x+y = 500
2nd: 0.09x+20 = .11y
Multiply 2nd by 100 to get:
3rd: 9x-11y=-2000
Multiply 1st by 9 to get:
4th: 9x+9y=4500
Subtract 3rd from 4th to get:
5th: 20y=2500
Then y=125
Substitute that into 1st to solve for "x", as follows:
x+125=500
x=375
Check this answer in 3rd:
9(375)-11(125)=-2000
2000=2000
It checks.
Answer: x=375; y=125
Cheers,
Stan H.